# TOWARDS AUDITORY ATTENTION DECODING WITH NOISE-TAGGING: A PILOT STUDY

### v1

@@
ABSTRACT: Auditory attention decoding (AAD) aims to extract from brain activity the attended speaker amidst candidate speakers, offering promising applications for neuro-steered hearing devices and brain-computer interfacing. This pilot study makes a first step towards AAD using the noise-tagging stimulus protocol, which evokes reliable code-modulated evoked potentials, but is minimally explored in the auditory modality. Participants were sequentially presented with two Dutch speech stimuli that were amplitude modulated with a unique binary pseudo-random noise-code, effectively tagging these with additional decodable information. We compared the decoding of unmodulated audio against audio modulated with various modulation depths, and a conventional AAD method against a standard method to decode noise-codes. Our pilot study revealed higher performances for the conventional method with 70 to 100 percent modulation depths compared to unmodulated audio. The noise-code decoder did not further improve these results. These fundamental insights highlight the potential of integrating noise-codes in speech to enhance auditory speaker detection when multiple speakers are presented simultaneously.
@@ INTRODUCTION
People suffering from hearing loss often have great difficulty in scenarios in which multiple individuals are speaking simultaneously, known as the ‘cocktail party scenario’, something which normal hearing persons have no difficulties with [1]. In these scenarios, hearing aids are not able to provide a good solution, as even though they are capable to suppress background noise, they are less capable of suppressing the unattended speakers. Some hearing aids attempt to mitigate this problem by using a heuristic, for example by enhancing the loudest or closest speaker, or the one who stands right in front of the listener. Unfortunately, these heuristics often lead to selecting the wrong speaker in real-life scenarios.
@@
As an alternative, auditory attention decoding (AAD) aims to decode the attended speaker from neural activity, as the synchronization between the listener’s brain signals and the attended speech envelope is stronger than with the ignored speech envelope [2]. This finding laid the groundwork for more research on hearing aids that allow for cognitive control, so called neuro-steered hearing aids [3]. These hearing aids aim to identify the attended speaker from neural activity, and correspondingly enhance this speaker’s audio signal whilst simultaneously suppressing the other speakers and background noise.
@@
The main idea behind such AAD approaches is to match the speech signals to the neural activity which synchronizes with the attended speech signal. For practical reasons typically electroencephalography (EEG) is used. Most AAD algorithms follow a stimulus reconstruction approach, also referred to as backward modeling or decoding. In this approach, the attended speech envelope is reconstructed from the EEG using a neural decoder, before the speaker whose envelope has the highest similarity with the reconstructed envelope is assumed to be the attended speaker [4]. Another approach is forward modeling or encoding, in which the objective is to predict the neural response from the speech envelopes via an encoder, and to compare these against the EEG [5, 6]. A third approach, sometimes referred to as the hybrid approach, combines decoding and encoding, by transforming both the speech envelopes and the EEG to minimize the irrelevant variance [7, 8]. Using such a hybrid approach, Geirnaert and colleagues [3] achieved a remarkable performance using canonical correlation analysis (CCA) to decode the attended speaker. Presenting audio from two simultaneous speakers, a mean accuracy of about 85 % was reached using 30 s decision windows. However, when decreasing the decision window length to about 10 s, the accuracy quickly dropped to below 80 %. This poses a significant limitation for real-world scenarios where fast speaker detection is crucial.
@@
Framing the speaker decoding problem as detecting which of several stimuli a person is attending to, another paradigm from the brain-computer interfacing (BCI) field recently reached remarkable performances. Specifically, the locus of visual attention can be decoded from EEG data using the code-modulated visual evoked potential (c-VEP). A c-VEP is the EEG response to pseudo-random visual stimulation sequences where stimuli are watermarked using noise-codes, a protocol called noise-tagging [9]. These noise-codes are selected or even optimized to be dissimilar, such that attending to one noise-code evokes substantially different brain activity than when attending to another code, facilitating the decoding of the attended stimulus. Such c-VEP BCIs have been reaching state-of-the-art performances up to 100 % classification accuracy using 1–4 s decision windows [10] or recently even within 300 ms [11] and a high number of stimuli, 29 and 40, respectively.
@@
In this pilot study, we make the first step towards auditory attention decoding using noise-tagging. Firstly, we aim to assess the feasibility of decoding the code-modulated auditory evoked potential (c-AEP), the response to auditory noise-tagging. Therefor, we use sequential presentation only, i.e., only one stimulus stimulus is presented at a time. Secondly, in this work we compare various modulation depths against no modulation. Thirdly, we compare how decoding approaches based on the speech envelope and noise-tag compare in terms of classification accuracy.
Successful implementation could make the step towards improving the decoding accuracy and speed in identifying the attended speaker. Furthermore, this exploration may pioneer a novel research avenue for the application of code-modulated responses in the auditory domain, a domain that has seen limited application compared to the visual modality, as so far only one study attempted this [12].
@@
Materials: The EEG data were recorded at a sample rate of 500 Hz with 64 active electrodes placed according to the 10-10 system and amplified by a BrainAmp (Brain Products GmbH) amplifier. The EEG data were preprocessed following a notch filter at 50 Hz, a bandpass filter between 1 and 20 Hz before resampling it to 120 Hz.
The auditory stimuli were two Dutch short stories [13], narrated by two different male speakers and recorded at 44100 Hz. They lasted approximately 6.5 min each and were a subset of the stimulus materials used by Das and colleagues [14]. Periods of silence exceeding 500 ms were truncated to 500 ms. The stimuli were normalized for loudness and presented dichotically to participants via headphones, with one story consistently presented to the left and the other to the right ear.
We used two 126-bit binary pseudo-random noise-codes from a family of modulated Gold codes [15, 16] to amplitude modulate the audio. These codes come in sets that are maximally uncorrelated with each other and each time-shifted versions of themselves. The codes were modulated to include only short (‘010’) and long (‘0110’) events. From the available modulated Gold codes, we carefully selected one that started with a 1, ended with a 0, and exhibited an almost uniform distribution of short and long events. The second code was a 61-bit phase-shifted version of the first. In this way, the noise-codes had identical properties, while minimizing auto-correlation at a maximum delay. The codes were presented at a bit rate of 40 Hz, and we always used the first code to modulate audio for the left ear, while the second code was always used to modulate audio for the right ear.
@@
Figure 1: Visualization of three different modulation depths using noise-tagging. Depicted is the unmodulated audio, i.e., 0 percent (blue), and 50 (gold) and 100 (brown) percent modulated audio. Additionally, shown are the smoothened noise-tags used for modulation (black). Audio was amplitude modulated by multiplying with the noise-code, retaining full audio amplitude when the code is 1, while only a percentage when it is zero. Therefore, the noise-code for 50 percent modulation ranges between 0.5–1, instead of 0–1 for 100 percent modulation. To ease comparison, we added the original audio (light gray) at the back of the modulated audio.
@@
Analysis: The classification of the attended speech in this work was done using two approaches, the conventional envelope CCA (further denoted as eCCA) in which only the information of the envelope of the speech signals is used, and reconvolution CCA (further denoted as rCCA), in which both the envelope as well as the noise-codes is used. Both approaches can be applied for the modulated audio and corresponding data, however only the eCCA approach can be applied on the unmodulated audio condition as here no code information is available for rCCA. The eCCA is based on the implementation by Geirnaert and colleagues [3]. The use and implementation of rCCA follows the work by Thielen and colleagues [10].
@@
Both methods make use of envelope information, which were obtained as follows: the modulated and unmodulated speech signals were each first filtered by a gammatone filterbank resulting in 15 frequency subbands for each speech stimulus [17]. Then, for each subband the absolute value was taken, followed by a power-law compression of 0.6. Next, each subband was lowpass filtered at 20 Hz and resampled to 120 Hz, which were a factor and multiple of the 40 Hz bitrate of the noise-code, respectively. Finally, for each stimulus all subbands were summed, with equal weights, to obtain the stimulus-specific envelopes.
@@
Instead of using the first component only, as in Eq. 2, using multiple components can improve classification accuracy but requires an additional classification model, e.g., a linear discriminant analysis (LDA) [3]. Specifically, CCA can deliver $K=\min(C,L)$ orthogonal components, ordered on decreasing canonical correlation. The $k$-th component contributes a spatial filter $\mathbf{w}_{k}$ and temporal filter $\mathbf{r}_{k}$, and delivers a Pearson’s correlation coefficient $\rho_{ki}$ following Eq. 2. These correlation coefficients across $K$ components (here $K=3$) are collected in a vector ${\bm{\rho}}_{i}$ for speaker $i$, and a feature vector $\mathbf{f}$ is created by subtracting the speakers’ canonical correlation vectors, $\mathbf{f}={\bm{\rho}}_{1}-{\bm{\rho}}_{2}$. The low-dimensional feature vector $\mathbf{f}$ can then be classified using a vanilla LDA, solving a binary classification problem of whether speaker 1 or speaker 2 was attended.
Reconvolution CCA: The rCCA approach consists of a template-matching classifier that predicts the attended speaker given the neural response evoked by the binary noise-code. The reconvolution model is based on the superposition hypothesis, stating that the response to a sequence of events is the linear summation of the responses evoked by the individual events [16].
@@
In this work, we extend the standard rCCA model from Thielen and colleagues [10] to incorporate envelope information. This is a crucial step, because a 1 in the code does not necessitate that there was audio in the stimulus. By incorporating the envelope and combining these with the events in the structure matrix, it can be avoided that an event is expected even though there was an audio amplitude of zero in the speech signal at that time. This is achieved by element-wise multiplying the event matrix $\mathbf{E}_{i}$ by the amplitudes of the envelope $\mathbf{a}_{i}$, before mapping the event matrix to a Toeplitz-like structure matrix.
@@
In this work, for rCCA, we only used the first CCA component for classification similar to the application of rCCA in the visual domain [10].
@@ DISCUSSION
This preliminary work aimed to contribute to our understanding of fundamental protocol design decisions when noise-tagging is integrated to perform AAD. We argue, that noise-tagging may provide additional information to the audio which could enhance AAD performance. In this pilot study with sequential presentation of the stimuli, we studied five different modulation conditions for the speech signals and compared two CCA-based decoding approaches. The results showed that the envelope CCA (eCCA) method achieved higher performances with full modulation than without modulation, while the alternative reconvolution CCA (rCCA) preferred a 70 percent modulation intensity to reach peak performance.
For all decision window lengths, the 100 and 70 modulation condition performed better than the unmodulated condition for the eCCA method. It could be speculated that this is due to the modulation that adds additional distinctive uncorrelated high-frequency content. Generally, a frequency range of 1-9 Hz is most informative for cortical tracking of speech envelopes [14]. Arguably, adding the noise-tags at a 40 Hz bitrate increases the frequency range of the envelope, which could increase the speed with which the envelopes become distinctive for eCCA. Therefor, we used a 20 Hz lowpass and a higher sampling frequency.
The lowest modulation depth that reached on par or higher performance with the unmodulated condition was a modulation of 70, both for using eCCA and rCCA. For both methods, a 50 modulation depth was able to reach similar performances as the 0 modulation, for decision window lengths of $\tau\leq 10$ s. Future work would need to assess the perception thresholds for modulation depths, in order to obtain least intrusive protocols for high usability.
@@
As this work is a preliminary study, a number of inherent limitations should be noted. First, the experiment only included five participants, of which two were authors with substantial experience as BCI users. This is a small sample size and needs to be enlarged in future studies. Second, in order to assess the feasibility of decoding a c-AEP response, this pilot used sequential stimulation instead of presenting the two speakers in parallel. Future studies need to investigate, if our observations generalize also to parallel stimulus presentation protocols.

### v2

@@
ABSTRACT: Auditory attention decoding (AAD) aims to extract from brain activity the attended speaker amidst candidate speakers, offering promising applications for neuro-steered hearing devices and brain-computer interfacing. This pilot study makes a first step towards AAD using the noise-tagging stimulus protocol, which evokes reliable code-modulated evoked potentials, but is minimally explored in the auditory modality. Participants were sequentially presented with two Dutch speech stimuli that were amplitude-modulated with a unique binary pseudo-random noise-code, effectively tagging these with additional decodable information. We compared the decoding of unmodulated audio against audio modulated with various modulation depths, and a conventional AAD method against a standard method to decode noise-codes. Our pilot study revealed higher performances for the conventional method with 70 to 100 percent modulation depths compared to unmodulated audio. The noise-code decoder did not further improve these results. These fundamental insights highlight the potential of integrating noise-codes in speech to enhance auditory speaker detection when multiple speakers are presented simultaneously.
@@ INTRODUCTION
People suffering from hearing loss often have great difficulty in scenarios in which multiple individuals are speaking simultaneously, known as the ‘cocktail party scenario’, something which normal hearing persons have no difficulties with [cherry1953]. In these scenarios, hearing aids are not able to provide a good solution, as even though they are capable to suppress background noise, they are less capable of suppressing the unattended speakers. Some hearing aids attempt to mitigate this problem by using a heuristic, for example by enhancing the loudest or closest speaker, or the one who stands right in front of the listener. Unfortunately, these heuristics often lead to selecting the wrong speaker in real-life scenarios.
@@
As an alternative, auditory attention decoding (AAD) aims to decode the attended speaker from neural activity, as the synchronization between the listener’s brain signals and the attended speech envelope is stronger than with the ignored speech envelope [ding2012]. This finding laid the groundwork for more research on hearing aids that allow for cognitive control, so called neuro-steered hearing aids [geirnaert2021]. These hearing aids aim to identify the attended speaker from neural activity, and correspondingly enhance this speaker’s audio signal whilst simultaneously suppressing the other speakers and background noise.
@@
The main idea behind such AAD approaches is to match the speech signals to the neural activity which synchronizes with the attended speech signal. For practical reasons typically electroencephalography (EEG) is used. Most AAD algorithms follow a stimulus reconstruction approach, also referred to as backward modeling or decoding. In this approach, the attended speech envelope is reconstructed from the EEG using a neural decoder, before the speaker whose envelope has the highest similarity with the reconstructed envelope is assumed to be the attended speaker [osullivan2015]. Another approach is forward modeling or encoding, in which the objective is to predict the neural response from the speech envelopes via an encoder, and to compare these against the EEG [lalor2010, ding2012b]. A third approach, sometimes referred to as the hybrid approach, combines decoding and encoding, by transforming both the speech envelopes and the EEG to minimize the irrelevant variance [cheveigne2018, dmochowski2018]. Using such a hybrid approach, Geirnaert and colleagues [geirnaert2021] achieved a remarkable performance using canonical correlation analysis (CCA) to decode the attended speaker. Presenting audio from two simultaneous speakers, a mean accuracy of about 85 % was reached using 30 s decision windows. However, when decreasing the decision window length to about 10 s, the accuracy quickly dropped to below 80 %. This poses a significant limitation for real-world scenarios where fast speaker detection is crucial.
@@
Framing the speaker decoding problem as detecting which of several stimuli a person is attending to, another paradigm from the brain-computer interfacing (BCI) field recently reached remarkable performances. Specifically, the locus of visual attention can be decoded from EEG data using the code-modulated visual evoked potential (c-VEP). A c-VEP is the EEG response to pseudo-random visual stimulation sequences where stimuli are watermarked using noise-codes, a protocol called noise-tagging [martinez2021]. These noise-codes are selected or even optimized to be dissimilar, such that attending to one noise-code evokes substantially different brain activity than when attending to another code, facilitating the decoding of the attended stimulus. Such c-VEP BCIs have been reaching state-of-the-art performances up to 100 % classification accuracy using 1–4 s decision windows [thielen2021] or recently even within 300 ms [shi2024] and a high number of stimuli, 29 and 40, respectively.
@@
In this pilot study, we make the first step towards auditory attention decoding using noise-tagging. Firstly, we aim to assess the feasibility of decoding the code-modulated auditory evoked potential (c-AEP), the response to auditory noise-tagging. Therefor, we use sequential presentation, i.e., only one stimulus is presented at a time. Secondly, in this work we compare various modulation depths against no modulation. Thirdly, we compare how decoding approaches based on the speech envelope and noise-tag compare in terms of classification accuracy.
Successful implementation could make the step towards improving the decoding accuracy and speed in identifying the attended speaker. Furthermore, this exploration may pioneer a novel research avenue for the application of code-modulated responses in the auditory domain, a domain that has seen limited application compared to the visual modality, as so far only one study attempted this [farquhar2008].
@@
Materials: The EEG data were recorded at a sample rate of 500 Hz with 64 active electrodes placed according to the 10-10 system and amplified by a BrainAmp (Brain Products GmbH) amplifier. The EEG data were preprocessed with a non-causal FIR notch filter at 50 Hz and a bandpass filterbetween 1 and 20 Hz before resampling to 120 Hz.For filtering, the default settings were used from the MNE toolbox, version 1.6.1 [Gramfort2013a].
The auditory stimuli were two Dutch short stories [radioboeken], narrated by two different male speakers and recorded at 44100 Hz. They lasted approximately 6.5 min each and were a subset of the stimulus materials used by Das and colleagues [das2016]. Periods of silence exceeding 500 ms were truncated to 500 ms. The stimuli were normalized for loudness and presented dichotically to participants via headphones, with one story consistently presented to the left and the other to the right ear.
We used two 126-bit binary pseudo-random noise-codes from a family of modulated Gold codes [gold1967, thielen2015] to amplitude-modulate the audio. The codes come in sets that are maximally uncorrelated with each other and each time-shifted versions of themselves. The codes were modulated to include only short (‘010’) and long (‘0110’) events. From the available modulated Gold codes, we carefully selected one that started with a 1, ended with a 0, and exhibited an almost uniform distribution of short and long events. The second code was a 61-bit phase-shifted version of the first. In this way, the noise-codes had identical properties, while minimizing auto-correlation at a maximum delay. The codes were presented at a bit rate of 40 Hz, and we always used the first code to modulate audio for the left ear, while the second code was always used to modulate audio for the right ear.
@@
Figure 1: Visualization of three different modulation depths using noise-tagging. Depicted is the unmodulated audio, i.e., 0 percent (blue), and 50 (gold) and 100 (brown) percent modulated audio. Additionally, shown are the smoothened noise-tags used for modulation (black). Audio was amplitude-modulated by multiplying with the noise-code, retaining full audio amplitude when the code is 1, while only a percentage when it is zero. Therefore, the noise-code for 50 percent modulation ranges between 0.5–1, instead of 0–1 for 100 percent modulation. To ease comparison, we added the original audio (light gray) at the back of the modulated audio.
@@
Analysis: The classification of the attended speech in this work was done using two approaches, the conventional envelope CCA (further denoted as eCCA) in which only the information of the envelope of the speech signals is used, and reconvolution CCA (further denoted as rCCA), in which both the envelope as well as the noise-codes is used. Both approaches can be applied for the modulated audio and corresponding data, however only the eCCA approach can be applied on the unmodulated audio condition as here no code information is available for rCCA. The eCCA is based on the implementation by Geirnaert and colleagues [geirnaert2021]. The use and implementation of rCCA follows the work by Thielen and colleagues [thielen2021].
@@
Both methods make use of envelope information, which were obtained as follows: the modulated and unmodulated speech signals were each first filtered by a gammatone filterbank resulting in 15 frequency subbands for each speech stimulus [biesmans2017]. Then, for each subband the absolute value was taken, followed by a power-law compression of 0.6. Next, each subband was lowpass filtered at 20 Hz and resampled to 120 Hz, which were a factor and multiple of the 40 Hz bitrate of the noise-code, respectively. Finally, for each stimulus all subbands were summed, with equal weights, to obtain the stimulus-specific envelopes.
@@
Instead of using the first component only, as in Eq. 2, using multiple components can improve classification accuracy but requires an additional classification model, e.g., a linear discriminant analysis (LDA) [geirnaert2021]. Specifically, CCA can deliver $K=\min(C,L)$ orthogonal components, ordered on decreasing canonical correlation. The $k$-th component contributes a spatial filter $\mathbf{w}_{k}$ and temporal filter $\mathbf{r}_{k}$, and delivers a Pearson’s correlation coefficient $\rho_{ki}$ following Eq. 2. These correlation coefficients across $K$ components (here $K=3$) are collected in a vector ${\bm{\rho}}_{i}$ for speaker $i$, and a feature vector $\mathbf{f}$ is created by subtracting the speakers’ canonical correlation vectors, $\mathbf{f}={\bm{\rho}}_{1}-{\bm{\rho}}_{2}$. The low-dimensional feature vector $\mathbf{f}$ can then be classified using a vanilla LDA, solving a binary classification problem of whether speaker 1 or speaker 2 was attended.
Reconvolution CCA: The rCCA approach consists of a template-matching classifier that predicts the attended speaker given the neural response evoked by the binary noise-code. The reconvolution model is based on the superposition hypothesis, stating that the response to a sequence of events is the linear summation of the responses evoked by the individual events [thielen2015].
@@
In this work, we extend the standard rCCA model from Thielen and colleagues [thielen2021] to incorporate envelope information. This is a crucial step, because a 1 in the code does not necessitate that there was audio in the stimulus. By incorporating the envelope and combining these with the events in the structure matrix, it can be avoided that an event is expected even though there was an audio amplitude of zero in the speech signal at that time. This is achieved by element-wise multiplying the event matrix $\mathbf{E}_{i}$ by the amplitudes of the envelope $\mathbf{A}_{i}$, before mapping the event matrix to a Toeplitz-like structure matrix.
@@
In this work, for rCCA, we only used the first CCA component for classification similar to the application of rCCA in the visual domain [thielen2021].
@@ DISCUSSION
This preliminary work aims to contribute to our understanding of fundamental protocol design decisions when noise-tagging is integrated to perform AAD. We argue that noise-tagging may provide additional information to the audio which could enhance AAD performance. In this pilot study with sequential presentation of the stimuli, we studied five modulation conditions for the speech signals and compared two CCA-based decoding approaches. The results showed that the envelope CCA (eCCA) method achieved higher performances with full modulation than without modulation, while the alternative reconvolution CCA (rCCA) preferred a 70 percent modulation intensity to reach peak performance.
For all decision window lengths, the 100 and 70 modulation conditions performed better than the unmodulated condition for eCCA. It could be speculated that this is due to the modulation adding distinctive uncorrelated high-frequency content. Generally, a 1-9 Hz range was found most informative for cortical tracking of speech envelopes [das2016]. Arguably, adding the noise-tags at a 40 Hz bitrate increases the envelope’s frequency range, which could increase the speed with which the envelopes become distinctive for eCCA. Therefore, we used a 20 Hz lowpass and a higher sampling frequency.
The lowest modulation depth that reached on par or higher performance with the unmodulated condition was a modulation of 70, both for using eCCA and rCCA. For both methods, a 50 modulation depth was able to reach similar performances as the 0 modulation, for decision window lengths of $\tau\leq 10$ s. Future work could assess the perception thresholds for modulation depths, to obtain the least intrusive protocols for high usability.
@@
As this work is a preliminary study, a number of inherent limitations should be noted. First, the experiment only included five participants, of which two were authors with substantial experience as BCI users. This is a small sample size, which included motivated participants, and needs to be enlarged in future studies. Second, in order to assess the feasibility of decoding a c-AEP response, this pilot used sequential stimulation instead of presenting the two speakers in parallel. Future studies need to investigate, if our observations generalize also to parallel stimulus presentation protocols.